# Preference-based Reinforcement Learning beyond Pairwise Comparisons: Benefits of Multiple Options

### v1

@@ H.2 Suboptimality Gap Under Sufficient Diversity Condition
$\lambda_{\min}\left(\Lambda_{T+1}\right)\geq\frac{\lambda_{0}}{4}(T-|{\mathcal{T}}^{w}|)K.$
Suppose $T-|{\mathcal{T}}^{w}|>0$. Then, combining the above results, we get
$\textup{{SubOpt}}(T)$ $\leq\beta_{T+1}(\delta)\frac{20\sqrt{2}}{\sqrt{\lambda_{0}(T-|{\mathcal{T}}^{w}|)K}}=\tilde{\mathcal{O}}\left(\sqrt{\frac{d}{\lambda_{0}(T-|{\mathcal{T}}^{w}|)K}}\right)$
@@
$=\tilde{\mathcal{O}}\left(\sqrt{\frac{d}{\lambda_{0}\big(T-\min\{(dK)^{2}/\kappa,T-1\}\big)K}}\right),$
where in the last inequality we use the fact that it holds that $|{\mathcal{T}}^{w}|\leq|{\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}|+|{\mathcal{T}}_{0}|=\tilde{\mathcal{O}}\left(\frac{d^{2}K^{2}}{\kappa}\right)$, which follows from Lemmas D.4 and D.5. This concludes the proof of Theorem H.1. ∎
H.3 Extension to Active Learning Setting
@@
This setting has received increasing attention in recent work [52, 49, 73, 19, 51, 79, 39]. However, most existing approaches focus exclusively on pairwise preference feedback. Mukherjee et al. [51] study an online learning-to-rank problem where, for each context, a fixed set of $K$ actions is provided, and the goal is to recover the true ranking based on feedback over these $K$ actions. In contrast, we consider a more general setting in which, for each context, a set of $N$ actions is available. The learner selects at most $K$ actions from this set and receives ranking feedback over the selected subset. Thekumparampil et al. [79] investigate the problem of ranking $N\geq K$ items using partial rankings over $K$ candidates, but under a context-free setting. In contrast, we study a stochastic contextual setting, where contexts are drawn from an unknown (and fixed) distribution.
In the active learning setting, the algorithm jointly selects the context $x_{t}$—which is no longer given but actively chosen—and the assortment $S_{t}$ by maximizing the average uncertainty objective. For computational efficiency, we employ the arbitrary reference action strategy described in Equation (H.1). (Note that one may alternatively use the reference action selection method from Equation (9), which selects $\bar{a}_{t}$ to maximize uncertainty.)
$(x_{t},S_{t})=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}}\mathop{\mathrm{argmax}}_{\begin{subarray}{c}S\in{\mathcal{S}}\\ \bar{a}_{t}\in S\end{subarray}}\frac{1}{|S|}\sum_{a\in S}\left\|\phi\left(x,a\right)-\phi\left(x,\bar{a}_{t}\right)\right\|_{H_{t}^{-1}},\quad\text{for any }\bar{a}_{t}\in\mathcal{A}.$ (H.3)
The rest of the algorithm proceeds in the same manner as Algorithm 3. With the above context-assortment selection strategy, M-AUPO achieves the following bound on the worst-case suboptimality gap, matching the order established in Theorem 1 (and in Theorem 2):
Theorem H.2.
Under the same setting as Theorem 1 and 2, with probability at least $1-\delta$, M-AUPO achieves the following worst-case suboptimality gap:
$\textup{{WorstSubOpt}}(T)=\begin{cases}\tilde{\mathcal{O}}\left(\frac{d}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}+\frac{d^{2}K^{2}}{\kappa T}\right),\quad\quad&\text{(PL loss)}\\[6.45831pt] \tilde{\mathcal{O}}\left(\frac{d}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}+\frac{d^{2}}{\kappa T}\right).\quad\quad&\text{(RB loss)}\end{cases}$
Proof of Theorem H.2.
We present only the proof using the PL loss (3), as extending it to the RB loss case (4) follows similarly to the extension from Theorem 1 to Theorem 2.
By the definition of the worst-cacse suboptimality gap, we have
@@
We adopt the same definitions for ${\mathcal{T}}_{0}$ (Equation (D.3)), ${\mathcal{T}}^{w}$ (Equation (D.1)), and $\Lambda_{t}$ (Equation (D.2)) as in Theorem 1. Then, we have
@@
$=\frac{1}{T}\sum_{t\in{\mathcal{T}}_{0}}\max_{x\in\mathcal{X}}\left[\big(\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)\right]$
$+\frac{1}{T}\sum_{t\in{\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}}\max_{x\in\mathcal{X}}\left[\big(\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)\right]$
$+\frac{1}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\max_{x\in\mathcal{X}}\left[\big(\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)\right]$
@@
$\leq\frac{8B}{\log(2)T}d\log\left(1+\frac{2K}{\log(2)\lambda}\right)+\frac{48\sqrt{2}BK^{2}}{\kappa T}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{2KT}{d\lambda}\right)$ (Lemma D.4 and D.5)
@@
$+\frac{1}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\max_{x\in\mathcal{X}}\left[\big(\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)\right].$ (H.4)
To further bound the last term of Equation (H.4), we get
@@
$\frac{1}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}$ $\max_{x\in\mathcal{X}}\left[\big(\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)\right]$
@@
$\leq\frac{1}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\max_{x\in\mathcal{X}}\left[\left\|\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\right\|_{H_{T+1}^{-1}}\left\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right\|_{H_{T+1}}\right]$ (Hölder’s ineq.)
@@
$\leq\frac{\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\max_{x\in\mathcal{X}}\left[\left\|\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\right\|_{H_{t}^{-1}}\right].$ ($H_{T+1}\succeq H_{t}$ and Corollary D.1, with prob. $1-\delta$)
Then, for any arbitrary $\bar{a}_{t}\in\mathcal{A}$, we have
$\frac{\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\max_{x\in\mathcal{X}}\left[\left\|\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\right\|_{H_{t}^{-1}}\right]$
$\leq\frac{\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\max_{x\in\mathcal{X}}\left[\left\|\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\bar{a}_{t}\right)\right\|_{H_{t}^{-1}}+\left\|\phi\left(x,\widehat{\pi}_{T}(x)\right)-\phi\left(x,\bar{a}_{t}\right)\right\|_{H_{t}^{-1}}\right]$
$\leq\frac{4\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\frac{1}{|S_{t}|}\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi(x_{t},\bar{a}_{t})\right\|_{H_{t}^{-1}}.$ ($(x_{t},S_{t})$ selection rule, Eqn. (H.3))
Hence, we further obtain
$\frac{4\beta_{T+1}(\delta)}{T}$ $\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\frac{1}{|S_{t}|}\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi(x_{t},\bar{a}_{t})\right\|_{H_{t}^{-1}}$
$\leq\frac{4\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\frac{1}{|S_{t}|}\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{H_{t}^{-1}}$ (Eqn. (H.3))
$\leq\frac{4\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\left(\frac{1}{|S_{t}|}\right)^{2}|S_{t}|}\sqrt{\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{H_{t}^{-1}}^{2}}$ (Cauchy-Schwartz ineq.)
$=\frac{4\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\left(\frac{1}{|S_{t}|}\right)^{2}|S_{t}|}\sqrt{50\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{\Lambda_{t}^{-1}}^{2}}$ (Lemma D.2, with prob. $1-\delta$)
$\leq\frac{30\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t=1}^{T}\min\left\{1,\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{\Lambda_{t}^{-1}}^{2}\right\}}$
$\leq\frac{30\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{2d\log\left(1+\frac{2KT}{d\lambda}\right)}$ (Lemma D.3)
$=\mathcal{O}\left(\frac{\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\cdot\sqrt{d\log\left(KT\right)}\right).$ (H.5)
Plugging Equation (H.5) into Equation (H.4), and setting $\beta_{T+1}(\delta)=\mathcal{O}\big(B\sqrt{d\log(KT)}+B\sqrt{\lambda}\big)$, with probability at least $1-3\delta$, we have
$\textup{{SubOpt}}(T)$ $=\tilde{\mathcal{O}}\left(\frac{d}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}+\frac{d^{2}K^{2}}{\kappa T}\right).$
Substituting $\delta\leftarrow\frac{\delta}{3}$, we conclude the proof of Theorem H.2. ∎
Appendix I Experimental Details and Additional Results
I.1 Synthetic Data
@@
Baselines. We evaluate our proposed algorithm, M-AUPO, against three baselines: (i) DopeWolfe [79], a method designed for non-contextual $K$-subset selection; (ii) Uniform, which selects assortments of size $K$ uniformly at random; and (iii) Best&Ref, which forms a pair of actions ($|S_{t}|=2$) by combining one action from the current policy with another from a reference policy (e.g., uniform random or SFT), following the setup in Online GSHF [89] and XPO [88].
Thekumparampil et al. [79] propose a D-optimal design approach for the Plackett-Luce objective to efficiently select informative subsets of items for comparison. Recognizing the computational complexity inherent in this method, they introduce a randomized Frank-Wolfe algorithm, named DopeWolfe, which approximates the optimal design by solving linear maximization sub-problems on randomly chosen variables. This approach reduces computational overhead while maintaining effective learning performance. However, their approach is specifically tailored to the single-context setting (e.g., Instance 2) and may not generalize well to the multiple-context scenarios (e.g., Instances 1, 3, and 4). While their original implementation updates the model parameters using a maximum likelihood estimation (MLE) procedure, we instead adopt an online update strategy (as described in Procedures 1 and 2) to ensure a fair comparison across all methods. For sampling size parameter $R$, we set $R=\min\{\binom{N}{K},100,000\}$.
The uniform random assortment selection strategy, Uniform, selects $K$ actions uniformly at random from the available action set $\mathcal{A}$ at each round, without utilizing any uncertainty or reward-based information. This approach can be effective when the feature representations are sufficiently diverse (e.g., Instances 1, 2, and 4), but may perform poorly when the diversity parameter $\lambda_{0}$ in Assumption H.1 is very small (e.g., Instance 3).
Best&Ref constructs an action pair ($|S_{t}|=2$) by combining two distinct sources of actions. The first action is chosen to maximize the current reward estimate, while the second is sampled from a reference policy—such as a uniform random policy or a supervised fine-tuned (SFT) model. This pairing mechanism follows the framework introduced in Online GSHF [89] and XPO [88]. In our experiments, we use the uniform random policy as the reference.
Performance measure. Since computing the exact suboptimality gap is challenging under a general distribution $\rho$, we instead evaluate the average realized regret, which serves as a slightly relaxed proxy for the suboptimality gap.
@@
$\leq\underbrace{\frac{1}{T}\sum_{t=1}^{T}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\pi_{t}(x_{t})\right)\big)^{\top}\bm{\theta}^{\star}}_{=:\textit{average realized regret}}+\tilde{\mathcal{O}}\left(\frac{1}{\sqrt{T}}\right),$
@@
Figure I.1: Performance comparisons for Instance 1 (Stochastic contexts) with $K=2$, 33, and $5$, evaluated under the PL loss (first row) and RB loss (second row).
Figure I.2: Performance comparisons for Instance 2 (Non-contextual) with $K=2$, 33, and $5$, evaluated under the PL loss (first row) and RB loss (second row).
Figure I.3: Performance comparisons for Instance 3 (Hard-to-learn contexts) with $K=2$, 33, and $5$, evaluated under the PL loss (first row) and RB loss (second row).
Figure I.4: Performance comparisons for Instance 4 (Skewed stochastic contexts) with $K=2$, 33, and $5$, evaluated under the PL loss (first row) and RB loss (second row).
Results. We present performance comparisons in Figures I.1 through I.4, corresponding to Instances 1 through 4, respectively. Overall, our algorithm, M-AUPO, consistently outperforms other baseline methods. The only exception is in Instance 2 (Figure I.2), a special case of the non-contextual setting, where M-AUPO performs slightly worse than DopeWolfe. This is an expected outcome, as DopeWolfe leverages a D-optimal design strategy, which is known to be highly effective in the single-context setting. However, it is important to note that DopeWolfe completely fails in more general contextual scenarios (Figures I.1, I.3, and I.4), and its computational cost is significantly higher than that of our approach (see Table I.1).
$K$ DopeWolfe Uniform Best&Ref M-AUPO (ours)
2 7.28 s 0.10 s 0.10s 1.94 s
3 99.6 s 0.18 s 0.10s 2.37 s
5 150.5 s 0.35 s 0.10s 2.94 s
7 218.8 s 0.58 s 0.10s 4.17 s
10 331.1 s 0.99 s 0.10s 4.50 s
Table I.1: Runtime comparison over 200 rounds (seconds)
The uniform random assortment selection strategy, Uniform, demonstrates competitive performance—though still worse than M-AUPO—in Instances 1, 2, and 4, as illustrated in Figures I.1, I.2, and I.4, respectively. However, in Instance 3 (Figure I.3), where the diversity parameter $\lambda_{0}$ is very small due to most feature vectors lying within a hyperplane, Uniform performs significantly worse, as discussed in Appendix H.2.
@@
Moreover, the suboptimality gap consistently decreases with larger $K$ across the three algorithms—M-AUPO, Uniform, and DopeWolfe—while Best&Ref shows no such improvement. For both M-AUPO and Uniform, this trend is consistent with our theoretical results (Theorems 1, 2, and H.1). In contrast, the improvement observed for DopeWolfe suggests that its current theoretical guarantees may be loose, as their bound actually deteriorates with increasing $K$ (recall that their theoretical guarantee worsens for larger $K$). This indicates that tighter bounds might be achievable by incorporating some of the techniques introduced in our work.
Table I.2 presents the average assortment size $|S_{t}|$ of M-AUPO for various values of the maximum assortment size $K$. In most cases, the algorithm selects the full $K$ actions, i.e., $|S_{t}|=K$. An exception occurs when $K$ is large (e.g., 30 or more), which may be impractical in real-world applications due to the increased annotation burden on human labelers.
@@
Table I.2: Assortment size $|S_{t}|$ of M-AUPO with varying maximum size $K$ in the synthetic data experiment
I.2 Real-World Dataset

### v2

@@ G.1 Avoiding $e^{B}$ Scaling in Regret Minimization
$\leq e(1+e)^{2}\sum_{t\notin{\mathcal{T}}^{w}}\frac{1}{|S^{\star}_{t}|}\sum_{a\in S^{\star}_{t}}\dot{\mu}\big(\psi_{t}(a,\widehat{a}_{t})^{\top}\widehat{\bm{\theta}}_{t}\big)\left\|\psi_{t}(a,\widehat{a}_{t})\right\|_{H_{t}^{-1}}$ ($S^{\star}_{t}:=\{\pi^{\star}(x_{t}),\widehat{a}_{t}\}$)
$\leq e(1+e)^{2}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}\frac{|S_{t}|}{|S^{\star}_{t}|}\sum_{a\in S^{\star}_{t}}\dot{\mu}\big(\psi_{t}(a,\widehat{a}_{t})^{\top}\widehat{\bm{\theta}}_{t}\big)\left\|\psi_{t}(a,\widehat{a}_{t})\right\|_{H_{t}^{-1}}^{2}}$ (Cauchy-Schwartz ineq., and ${\mathcal{T}}^{w}\subseteq[T]$)
$\leq e(1+e)^{2}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}\frac{\cancel{|S_{t}|}}{\cancel{|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}S_{t}}|}}\sum_{a\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}S_{t}}}\dot{\mu}\big(\psi_{t}(a,\widehat{a}_{t})^{\top}\widehat{\bm{\theta}}_{t}\big)\left\|\psi_{t}(a,\widehat{a}_{t})\right\|_{H_{t}^{-1}}^{2}}$
$\leq\sqrt{2}e^{2}(1+e)^{2}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}\sum_{a\in S_{t}}\dot{\mu}\big(\psi_{t}(a,\widehat{a}_{t})^{\top}\widehat{\bm{\theta}}_{t}\big)\left\|\psi_{t}(a,\widehat{a}_{t})\right\|_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Sigma_{t}^{-1}}}^{2}}$ (Lemma D.1)
$=\mathcal{O}\!\left(\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\cdot\sqrt{d\log\left(KT\right)}\right),$ (Lemma E.6)
where, in the last inequality, we introduce $\Sigma_{t}=\!\!\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\!\sum_{a\in S_{s}}\dot{\mu}\left(\psi_{t}(a,\widehat{a}_{t})^{\top}\widehat{\bm{\theta}}_{s}\right)\psi_{t}(a,\widehat{a}_{t})\psi_{t}(a,\widehat{a}_{t})^{\top}+\lambda\mathbf{I}_{d},$ and use the relation $H_{t}\succeq\frac{1}{2e^{2}}\Lambda_{t}\succeq\frac{1}{2e^{2}}\Sigma_{t}$ ($\Lambda_{t}$ is defined in Equation (D.1)). In the final equality, we apply the fact that if $\lambda=d\log T$ and $T\geq e^{K/d}$, then, for any $t$,
$\dot{\mu}\big(\psi_{t}(a,\widehat{a}_{t})^{\top}\widehat{\bm{\theta}}_{t}\big)\left\|\psi_{t}(a,\widehat{a}_{t})\right\|_{\Sigma_{t}^{-1}}^{2}\leq\frac{2K}{\lambda}\leq 2.$
Therefore, using $\beta_{T}(\delta)=\mathcal{O}\big(B\sqrt{d\log(KT)}+B\sqrt{\lambda}\big)$, we obtain, with probability at least $1-\delta$,
@@
$\textup{{Reg}}_{T}=\tilde{\mathcal{O}}\left(d\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}+\frac{d^{2}}{\kappa}\right).$
This completes the proof of Theorem G.1. ∎
G.2 Extension to Active Learning Setting
@@
This setting has received increasing attention in recent work [muldrew2024active, mehta2023sample, scheid2024optimal, das2024active, mukherjee2024optimal, thekumparampil2024comparing, kveton2025active]. However, most existing approaches focus exclusively on pairwise preference feedback. mukherjee2024optimal study an online learning-to-rank problem where, for each context, a fixed set of $K$ actions is provided, and the goal is to recover the true ranking based on feedback over these $K$ actions. In contrast, we consider a more general setting in which, for each context, a set of $N$ actions is available. The learner selects at most $K$ actions from this set and receives ranking feedback over the selected subset. thekumparampil2024comparing investigate the problem of ranking $N\geq K$ items using partial rankings over $K$ candidates, but under a context-free setting. In contrast, we study a stochastic contextual setting, where contexts are drawn from an unknown (and fixed) distribution.
In the active learning setting, the algorithm jointly selects the context $x_{t}$—which is no longer given but actively chosen—and the assortment $S_{t}$ by maximizing the average uncertainty objective. For each candidate context $x$, it first constructs the assortment $S_{t}(x)$ by solving Equation (11). It then selects $x_{t}=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}}f_{t}(S_{t}(x))$, and sets $S_{t}=S_{t}(x_{t})$. The rest of the algorithm proceeds in the same manner as Algorithm 3.
For simplicity, we focus on the RB loss and its corresponding update rule (Procedure 2). Then, we can obtain the following bound on the worst-case suboptimality gap.
Theorem G.2.
Under the same setting as Theorem 2, with probability at least $1-\delta$, the selection rule for $(x_{t},S_{t})$ described above achieves:
$\textup{{WorstSubOpt}}(T)=\tilde{\mathcal{O}}\left(\frac{d}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}+\frac{d^{2}}{\kappa T}\right).$
Proof of Theorem G.2.
By the definition of the worst-case suboptimality gap, we have
@@
We adopt the same definitions for ${\mathcal{T}}^{w}:=\{t\in[T]:\max_{a,a^{\prime}\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|_{H_{t}^{-1}}\geq 1/\beta_{T+1}(\delta)\}$. Thus, we get
@@
$\leq\mathcal{O}\left(\frac{B}{\kappa T}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{T}{d\lambda}\right)\right)$ (Lemma D.2)
@@
$+\frac{1}{T}\sum_{t\notin{\mathcal{T}}^{w}}\max_{x\in\mathcal{X}}\left[\big(\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)\right].$ (G.2)
To further bound the last term of Equation (G.2), we get
@@
$\frac{1}{T}$ $\sum_{t\notin{\mathcal{T}}^{w}}\max_{x\in\mathcal{X}}\left[\big(\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)\right]$
@@
$\leq\frac{1}{T}\sum_{t\notin{\mathcal{T}}^{w}}\max_{x\in\mathcal{X}}\left[\left\|\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\right\|_{H_{T+1}^{-1}}\right]\left\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right\|_{H_{T+1}}$ (Hölder’s ineq.)
@@
$\leq\frac{\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}^{w}}\max_{x\in\mathcal{X}}\left[\left\|\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\right\|_{H_{t}^{-1}}\right].$ ($H_{T+1}\succeq H_{t}$ and Corollary C.1, with prob. $1-\delta$)
For simplicity, let $x_{t}^{\star}=\mathop{\mathrm{argmax}}_{x}\left\|\phi\left(x,\pi^{\star}(x)\right)-\phi\left(x,\widehat{\pi}_{T}(x)\right)\right\|_{H_{t}^{-1}}$. It therefore suffices to bound $\left\|\phi\left(x_{t}^{\star},\pi^{\star}(x)\right)-\phi\left(x_{t}^{\star},\widehat{\pi}_{T}(x)\right)\right\|_{H_{t}^{-1}}$. From here, the same argument as in the proof of Theorem 2 applies, with the only modification that the procedure now selects both $x_{t}$ and $S_{t}$. This completes the proof of Theorem G.2. ∎
Figure G.1: Performance comparisons for Instance 1 (Stochastic contexts) with $K=2$, 33, and $5$, evaluated under the PL loss (first row) and RB loss (second row).
Appendix H Experimental Details and Additional Results
H.1 Synthetic Data
@@
Baselines. We evaluate our proposed algorithm, M-AUPO, against three baselines: (i) DopeWolfe [thekumparampil2024comparing], a method designed for non-contextual $K$-subset selection; (ii) Uniform, which selects assortments of size $K$ uniformly at random; and (iii) Best&Ref, which forms a pair of actions ($|S_{t}|=2$) by combining one action from the current policy with another from a reference policy (e.g., uniform random or SFT), following the setup in Online GSHF [xiong2023iterative] and XPO [xie2024exploratory].
When using the PL loss–based update in our algorithm, M-AUPO, the exact expectation over rankings is computationally expensive. To mitigate this, we approximate the expectation via Monte Carlo sampling with $5$ samples. Consequently, the computational cost of the PL-based approach is higher than that of the RB-based approach (see Table H.1).
thekumparampil2024comparing propose a D-optimal design approach for the Plackett-Luce objective to efficiently select informative subsets of items for comparison. Recognizing the computational complexity inherent in this method, they introduce a randomized Frank-Wolfe algorithm, named DopeWolfe, which approximates the optimal design by solving linear maximization sub-problems on randomly chosen variables. This approach reduces computational overhead while maintaining effective learning performance. However, their approach is specifically tailored to the single-context setting (e.g., Instance 2) and may not generalize well to the multiple-context scenarios (e.g., Instances 1, 3, and 4). While their original implementation updates the model parameters using a maximum likelihood estimation (MLE) procedure, we instead adopt an online update strategy (as described in Procedures 1 and 2) to ensure a fair comparison across all methods. For sampling size parameter $R$, we set $R=\min\{\binom{N}{K},100,000\}$.
The uniform random selection strategy, Uniform, selects $K$ actions uniformly at random from the available action set $\mathcal{A}$ at each round, without utilizing any uncertainty or reward-based information.
Best&Ref constructs an action pair ($|S_{t}|=2$) by combining two distinct sources of actions. The first action is chosen to maximize the current reward estimate, while the second is sampled from a reference policy—such as a uniform random policy or a supervised fine-tuned (SFT) model. This pairing mechanism follows the framework introduced in Online GSHF [xiong2023iterative] and XPO [xie2024exploratory]. In our experiments, we use the uniform random policy as the reference.
Performance measure. Since computing the exact suboptimality gap is challenging under a general distribution $\rho$, we instead evaluate the realized regret, which serves as a slightly relaxed proxy for the suboptimality gap.
@@
$\leq\underbrace{\frac{1}{T}\sum_{t=1}^{T}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\pi_{t}(x_{t})\right)\big)^{\top}\bm{\theta}^{\star}}_{=:\textit{realized regret}}+\tilde{\mathcal{O}}\left(\frac{1}{\sqrt{T}}\right),$
@@
Figure H.1: Performance comparisons for Instance 2 (Non-contextual) with $K=2$, 33, and $5$, evaluated under the PL loss (first row) and RB loss (second row).
Figure H.2: Performance comparisons for Instance 3 (Hard-to-learn contexts) with $K=2$, 33, and $5$, evaluated under the PL loss (first row) and RB loss (second row).
Figure H.3: Performance comparisons for Instance 4 (skewed stochastic contexts) with $K=2$, 33, and $5$, evaluated under the PL loss (first row) and RB loss (second row).
Results. We present performance comparisons in Figures G.1 through H.3, corresponding to Instances 1 through 4, respectively. Overall, our algorithm, M-AUPO, consistently outperforms other baseline methods. The only exception is in Instance 2 (Figure H.1), a special case of the non-contextual setting, where M-AUPO performs slightly worse than DopeWolfe. This is an expected outcome, as DopeWolfe leverages a D-optimal design strategy, which is known to be highly effective in the single-context setting. However, it is important to note that DopeWolfe completely fails in more general contextual scenarios (Figures G.1, H.2, and H.3), and its computational cost is significantly higher than that of our approach (see Table H.1).
$K$ DopeWolfe Uniform Best&Ref M-AUPO (PL) M-AUPO (RB)
2 7.28 s 0.10 s 0.10s 1.09 s 1.04 s
3 99.6 s 0.18 s 0.10s 2.35 s 1.06 s
5 150.5 s 0.35 s 0.10s 2.71 s 1.10 s
7 218.8 s 0.58 s 0.10s 4.10 s 1.13 s
10 331.1 s 0.99 s 0.10s 6.89 s 1.31 s
Table H.1: Runtime comparison over 200 rounds (seconds)
The uniform random assortment selection strategy, Uniform, shows reasonable performance—though consistently inferior to M-AUPO—in Instances 1, 2, and 4, as presented in Figures G.1, H.1, and H.3, respectively. In contrast, for Instance 3 (Figure H.2), where most features are uninformative (being nearly orthogonal to the true parameter) and thus careful action selection becomes crucial, Uniform performs substantially worse than M-AUPO.
@@
Moreover, the suboptimality gap consistently decreases with larger $K$ across the three algorithms—M-AUPO, Uniform, and DopeWolfe—whereas Best&Ref shows no such improvement, as it always selects only two actions regardless of $K$. For M-AUPO, this trend is consistent with our theoretical results (Theorems 1 and 2). In contrast, the improvement observed for DopeWolfe suggests that its current theoretical guarantees may be loose, as their bound actually deteriorates with increasing $K$ (recall that their theoretical guarantee worsens for larger $K$). This suggests that tighter bounds could be obtained by incorporating techniques similar to those introduced in our work.
Table H.2 presents the average assortment size $|S_{t}|$ of M-AUPO for various values of the maximum assortment size $K$. In most cases, the algorithm selects the full $K$ actions, i.e., $|S_{t}|=K$. An exception occurs when $K$ is large (e.g., 30 or more), which may be impractical in real-world applications due to the increased annotation burden on human labelers.
@@
Table H.2: Assortment size $|S_{t}|$ of M-AUPO with varying $K$ in the synthetic experiment
H.2 Real-World Dataset